OpenStudy (anonymous):

Problem: My friend sets out walking at a speed of 3 miles per hour. I set out behind her 5 minutes later at 4 miles per hour. For how many minutes will my friend have been walking when I catch up to her?

OpenStudy (anonymous):

pls answer

OpenStudy (danjs):

distance = rate * time i would use a universal time for both, and call hers T and yours T-5 since you add no distance for the first 5 min when you catch up to her, both distances are the same, and equate them 3 * T = 4 * (T - 5) she moves at 3mph for T time, and you move at 4mph for T-5 time solve for Time T when you catch up

OpenStudy (danjs):

i just realized T is in minutes and the rate is in hour, since they want minutes , change the rates to miles per minute (3/60) * T = (4/60) * (T - 5)